# Can Input Attributions Interpret the Inductive Reasoning Process in In-Context Learning?

### v2

@@ Abstract
Interpreting the internal process of neural models has long been a challenge. This challenge remains relevant in the era of large language models (LLMs) and in-context learning (ICL); for example, ICL poses a new issue of interpreting which example in the few-shot examples contributed to identifying/solving the task. To this end, in this paper, we design synthetic diagnostic tasks of inductive reasoning, inspired by the generalization tests in linguistics; here, most in-context examples are ambiguous w.r.t. their underlying rule, and one critical example disambiguates the task demonstrated. The question is whether conventional input attribution (IA) methods can track such a reasoning process, i.e., identify the influential example, in ICL. Our experiments provide several practical findings; for example, a certain simple IA method works the best, and the larger the model, the generally harder it is to interpret the ICL with gradient-based IA methods.
@@
Can Input Attributions Interpret
the Inductive Reasoning Process in In-Context Learning?
Mengyu Ye1     Tatsuki Kuribayashi2     Goro Kobayashi1,3     Jun Suzuki1 1Tohoku University     2MBZUAI     3RIKEN {ye.mengyu.s1, goro.koba}@dc.tohoku.ac.jp tatsuki.kuribayashi@mbzuai.ac.ae     jun.suzuki@tohoku.ac.jp
@@ 1 Introduction
In past years, input attribution (IA) methods, e.g., gradient norm Simonyan et al. (2014); Li et al. (2016a), have typically been employed in the natural language processing (NLP) field to interpret input–output associations exploited by neural NLP models Vinyals and Le (2015); Li et al. (2016b). Recently, large language models (LLMs) and mechanistic interpretability (MI) research Olah et al. (2020); Bereska and Gavves (2024) have shifted the research focus to understanding the circuits within LLMs by intervening in their internal representations. Despite the enriched scope of research, such rapid progress has missed some intriguing questions bridging the IA and MI eras: in particular, do conventional IA methods still empirically work in the modern NLP setting, specifically in the context of LLMs and in-context learning (ICL)?
In this paper, we revisit IA methods in interpreting LLM-based ICL Brown et al. (2020). Specifically, we assess how well IA methods can track the most influential example in a few-shot examples. This question is worth investigating for several reasons. First, input attribution would still serve as a necessary and sufficient explanation in typical practical cases; some users might simply seek which part of the context is heavily referred to by an LLM system rather than LLMs’ internal processes identified by MI methods. Second, the modern NLP setting, specifically ICL, differs from the conventional settings where IA methods have been tested — identifying the input-output association within a specific test instance ($X_{k}$, $y_{k}$). In contrast, applying IA methods to the entire ICL input (few-shot examples) already entails tracking the learning process as well as the input-output association within a specific target instance. That is, this extended scope includes the interpretation of which example among the demonstrations [$(X_{1},y_{1}),\cdots,(X_{k-1},y_{k-1})$] contributed to identifying the targeted task/rule and then answer a target question $X_{k}$. This is rather a type of instance-based interpretation of neural models Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021), and it has been little explored such interpretation is feasible with IA methods.
@@
Our experiments reveal several findings:
  • Gradient norm, the simplest IA method, frequently outperforms other interpretability methods (e.g., integrated gradient), suggesting that the advantage of more recently proposed IA methods does not always generalize in interpreting ICL×$\times$×LLM.
  • Our tested interpretability methods, including simple gradient norm, did not work stably across different tasks and models, posing their general limitations in interpreting ICL with IA methods.
  • Different interpretability methods exhibited different properties with respect to scaling the number of in-context examples or model size; for example, IA methods perform better in many-shot scenarios, whereas a particular baseline interpretability method (i.e., self-answer) works well on larger models.
@@ 2.1 Input attribution (IA) methods
Input attribution (IA) methods are commonly-used techniques for interpreting and explaining the predictions of machine learning models (Denil et al., 2014; Li et al., 2016a; Poerner et al., 2018; Arras et al., 2019, etc.). Specifically, IA methods determine how much each input feature contributes to a particular prediction; that is, given input tokens $X\coloneqq[x_{1},\ldots,x_{n}]$ and output $y$, the IA methods yield the strength of contribution $S(x_{i})$ of each input $x_{i}$ to the output $y$. Note that the input $X$ in ICL consists of several in-context examples (§ 2.2), and the answer to the target question is denoted as $y$. We examine the following four representative IA methods in the ICL×$\times$×LLM context:
@@
For the IE, GN, and I×$\times$×G methods, we adopt a contrastive explanation setting, which Yin and Neubig (2022) have shown to be quantitatively superior to the original non-contrastive setting. IA methods in this setting measure how much an input token $x_{i}$ influences the model to increase the probability of target token $y_{t}$ while decreasing that of foil token $y_{f}$. A foil token can be defined as an output with an alternative, incorrect generalization (§ 3). Contrastive versions of IE, GN, and I×$\times$×G are defined as follows:
@@ 2.2 Interpreting in-context learning (ICL)
We focus on the ICL setting Brown et al. (2020), which has typically been adopted in modern LLM-based reasoning. An input prompt in ICL setting consists of few-shot examples $E$ and a target question. $E$ is composed of $n$ examples [$e_{1},\cdots,e_{n}$], each of which contains an input-output pair $e_{i}=(X_{i},f(X_{i}))$, given a function $f$ associated with the task. Let $X_{n+1}$ represent the target question $q$ that the model must answer. The ICL setting is formed as follows:
$\underbrace{\overbrace{\overbrace{X_{1},f(X_{1})}^{\textit{example}~{}e_{1}},\ldots,\overbrace{X_{n},f(X_{n})}^{\textit{example}~{}e_{n}}}^{\textit{few-shot examples E}},\overbrace{X_{n+1}}^{q}}_{\textit{prompt}},\underbrace{f(X_{n+1})}_{\textit{completion}}$
@@
Notably, we consider the use of IA methods to identify a particular example $e^{*}\in E$ in input, instead of a particular token. To compute an IA score for an example $S(e_{i})$, we sum up the IA scores for its constituent tokens: $S(e_{i})=\sum_{x_{j}\in(X_{i},y_{i})=e_{i}}S(x_{j})$. [^2] An exception applies in the IE method; the attribution score for an example $e_{i}$ is simply computed by erasing the corresponding $X_{i}$ and $f(X_{i})$ from the input sequence. Our interest is which example obtains the highest IA score, i.e., $\operatorname*{arg\,max}_{e_{i}\in E}S(e_{i})$.
@@ 3 Problem settings
Figure 2: Attribution accuracies for each task/model in the 10-shot setting (thus, the chance rate is 20% and 10% for the top-two and top-one metric, respectively; red dotted line). The edit distance and attention baselines are indicated by a black dotted line and gray bar, respectively.
In addition to them, we adopted a simple task of associative recall (AR), which is typically employed in studying ICL, where the model is supposed to simply memorize the key: value mapping rules demonstrated in the prompt and apply them to the target question. Linguists may be more interested in the task of, for example, syntactic transformation to an interrogative sentence McCoy et al. (2020) based on the original poverty of the stimulus argument in the language domain Chomsky (1980). However, such a realistic task interferes with the models’ meta-linguistic knowledge; thus, we adopted artificial (and somewhat simpler) ones.
@@ 4.1 Overview
For each synthetic task, we create 360 different questions with different sets of few-shot examples and a target question. In the LD, AM, VO, TA, and PT tasks, the correct rule is selected out of the two candidates (rules A or B shown in Table 1) in a 1:1 ratio. The position of the most influential (i.e., disambiguating) example $e^{*}$ is assigned according to a uniform distribution over all positions except the first. We test IA methods using only the questions that models answered correctly.
@@
We evaluate five LLMs: Llama-2-7B, Llama-2-13B Touvron et al. (2023), Gemma-2-2B, Gemma-2-9B, and Gemma-2-27B Riviere et al. (2024). As a prerequisite of our experiments, the models should be able to learn the task, i.e., be sufficiently sensitive to the disambiguating example $e^{*}$ and use this to determine the correct rule. To ensure this ability, we fine-tune these models on each task (see Appendix A), but the conclusions overall did not alter before and after fine-tuning (see Appendix B).
@@ 4.3 Baselines
We also examine directly asking the models to generate their rationale. Specifically, we have models generate the most informative example in a prompt (Appendix D) in deriving their answer as a post-hoc explanation. This might be, more or less, relevant to the verbalization of aha moment recently observed in DeepSeek models Guo et al. (2025).
@@ 5 Experiments
Figure 2 shows the results in 10-shot settings, with both top-2 and top-1 metrics. Additional analyses are presented in Appendices CE, and F.
@@ 5.2 Scaling properties
Figure 3: Attribution accuracy for interpreting Gemma-2-2B models across all six tasks. Gradient-based methods are relatively robust to the number of few-shot examples, while there is a consistent, large drop in attribution accuracy in SA. Note that both x-axis and y-axis are in log scale.
@@
Next, given the trend of long-context LLMs, we examine the relationship between attribution accuracy and the number of few-shot examples. Figure 3 shows the attribution accuracy for Gemma-2-2B across all six tasks in different number of in-context examples. This demonstrates that gradient-based methods maintain accuracy or rather improve against the longer context, in contrast to the decreasing chance rate. This suggests the robustness of IA methods in long-context scenarios, highlighting their potential for interpreting inputs with extensive contextual information. Notably, the quality of self-answer consistently degraded as the number of in-context examples increased, while the positive scaling effect was observed in the previous analysis of model size. That is, the gradient-based methods and self-answer approach exhibit an insightful trade-off between different scaling properties.
@@
SP (Rule A) $98.0$ $52.5$ $50.0$
SP (Rule B) $2.0$
ST (Rule A) $68.5$ $59.0$ $51.1$
ST (Rule B) $31.5$
@@ 6 Discussion
Our tasks can not be answered without disambiguating aha examples. Thus, it is somewhat unintuitive to see the non-100% attribution accuracy of the IE method (again, the LLMs understood the task as they achieved 100% accuracy in the tasks) — what happens here? To obtain a hint to clarify IE’s potential limitations, we analyze model behaviors when the disambiguating example is excluded. Interestingly, LLMs adopted a specific generalization (rule) in each task when there was no disambiguating example (Table 2); in other words, they sometimes exhibited strong inductive biases in our tasks. That is, when the correct rule is equal to their preferred rule by their inductive bias, they can answer the task correctly even without disambiguating examples, and the IE method does not compute a proper attribution score. It is now common to see LLMs have particular inductive biases (not a tabula rasaWarstadt et al. (2020b); Kharitonov and Chaabouni (2021). Catching such generalization bias with IA methods represents an inherent challenge, highlighting their potential limitations in interpreting ICL.
Why was I×$\times$×G worse than gradient norm?
@@
One advantage of the I×$\times$×G method, compared to the simple gradient norm, is the consideration of the norm of the input embedding Shrikumar et al. (2017); Denil et al. (2014). Since a large vector tends to have a large dot product with another vector, the norm of the input token embedding (vector) is expected to affect the IA score of I×$\times$×G. Then, no improvement of I×$\times$×G over the simple gradient norm suggests that, at least in our settings, the norm of the embeddings was not informative to estimate the input attribution. The norm of the embedding largely has decontextualized information about the word, such as frequency, and it may make sense that such information is not helpful to interpreting our controlled, artificial ICL tasks consisting of alphabet characters, numbers, or random words.
Why was IG worse than gradient norm?
@@
IG is a path-based approach; the gradient is accumulated from a baseline vector (typically zero vector) to the targeted input representation (in our case, the sequence of input embeddings representing few-shot examples). This approach is somewhat intuitive when considering an attribution for a particular word or sentence; for example, suppose one computes an attribution to the word “excellent” in an input for a particular task-specific model, IG may trace the path from zero to the “excellent” vector, which will be in a kind of the goodness direction, involving the points corresponding to, e.g., “okay” “decent,” “good,” “excellent” Sanyal and Ren (2021). Then, one critical question is — what does this path mean in the prompt/task space? Different prompt representations will no longer correspond to the same task; thus, the attribution of a particular token in the middle of such a path in the prompt space may no longer be an attribution under a targeted task. This can be one concern toward the ineffectiveness of IG in our settings.

### v5

# Can Input Attributions Explain Inductive Reasoning in In-Context Learning?

@@ Abstract
Interpreting the internal process of neural models has long been a challenge. This challenge remains relevant in the era of large language models (LLMs) and in-context learning (ICL); for example, ICL poses a new issue of interpreting which example in the few-shot examples contributed to identifying/solving the task. To this end, in this paper, we design synthetic diagnostic tasks of inductive reasoning, inspired by the generalization tests typically adopted in psycholinguistics. Here, most in-context examples are ambiguous w.r.t. their underlying rule, and one critical example disambiguates it. The question is whether conventional input attribution (IA) methods can track such a reasoning process, i.e., identify the influential example, in ICL. Our experiments provide several practical findings; for example, a certain simple IA method works the best, and the larger the model, the generally harder it is to interpret the ICL with gradient-based IA methods.
@@
Can Input Attributions Explain Inductive Reasoning
in In-Context Learning?
Mengyu Ye1     Tatsuki Kuribayashi2,1     Goro Kobayashithanks: Currently affiliated with Preferred Networks, Inc.1,3     Jun Suzuki1,3 1Tohoku University     2MBZUAI     3RIKEN {ye.mengyu.s1, goro.koba}@dc.tohoku.ac.jp tatsuki.kuribayashi@mbzuai.ac.ae     jun.suzuki@tohoku.ac.jp
@@ 1 Introduction
In the natural language processing (NLP) field, input attribution (IA) methods, e.g., relying on gradient norm Simonyan et al. (2014); Li et al. (2016a), have typically been employed to interpret input–output associations exploited by neural NLP models Vinyals and Le (2015); Li et al. (2016b). Recently, as large language models (LLMs) and mechanistic interpretability (MI) research Olah et al. (2020); Bereska and Gavves (2024) has gained attention, the research focus is leaning toward understanding the circuits within LLMs by intervening in their internal representations and information flow. Despite the enriched scope of research, such rapid progress has missed some intriguing questions bridging the IA and MI eras: in particular, do conventional IA methods still empirically work in the modern NLP setting, specifically with LLMs and in-context learning (ICL)?
In this paper, we revisit IA methods in interpreting LLM-based ICL Brown et al. (2020). Specifically, we assess how well IA methods can track the most influential example in a few-shot examples. This question is worth investigating for several reasons. First, input attribution would still serve as a necessary and sufficient explanation in typical practical cases; some users might simply seek which part of the context is heavily referred to by an LLM system rather than LLMs’ internal processes identified by MI methods. Second, the modern NLP setting, specifically ICL, differs from the conventional settings where IA methods have been tested — identifying the input-output association within a specific instance ($X_{k}$, $y_{k}$). In contrast, applying IA methods to the entire ICL input (few-shot examples) entails tracking the model’s learning process over these instances as well as the pure input-output association within a specific target instance. This involves the interpretation of which example among the demonstrations [$(X_{1},y_{1}),\cdots,(X_{k-1},y_{k-1})$] contributed to identifying the targeted task/rule and then answering a target question $X_{k}$. This is rather a type of instance-based interpretation of neural models Wachter et al. (2017); Charpiat et al. (2019); Hanawa et al. (2021), and it has been little explored such interpretation is feasible with IA methods.
@@
In our experiments, IA methods are compared to other common interpretability approaches relying on, such as attention weights and post-hoc explanation generations. The experimental results demonstrate:
  • Gradient norm, the simplest IA method, frequently outperforms other interpretability methods (e.g., integrated gradient, attention weights), suggesting that the advantage of more recently proposed IA methods does not always generalize in interpreting ICL×$\times$×LLM.
  • We revealed that different interpretability methods exhibited different advantages in interpreting the ICL process, against scaling in several aspects. For example, IA methods perform better in many-shot scenarios, whereas a post-hoc explanation generations work well on larger models.
  • Our tested interpretability methods, including simple gradient norm, did not work stably across different tasks and models, posing their general limitations in interpreting ICL with IA methods. Some existing IA methods frequently failed even on a very simple associative recall task, and there is room to sophisticate previously developed interpretability tools to be suitable for LLMs.
@@ 2.1 Input attribution (IA) methods
Input attribution (IA) methods are commonly-used techniques for interpreting and explaining the predictions of machine learning models (Denil et al., 2014; Li et al., 2016a; Poerner et al., 2018; Arras et al., 2019, etc.). Specifically, IA methods determine how much each input feature contributes to a particular prediction; that is, given input tokens $X\coloneqq[x_{1},\ldots,x_{n}]$ and output $y$, the IA methods yield the strength of contribution $S(x_{i})$ of each input token $x_{i}\in X$ to the output $y$. Note that the input $X$ in ICL consists of tokens of several in-context examples (§ 2.2), and the answer to the target question is denoted as $y$. We examine the following four representative IA methods in the ICL×$\times$×LLM context:
@@
For the IE, GN, and I×$\times$×G methods, we adopt a contrastive explanation setting, which Yin and Neubig (2022) have shown to be quantitatively superior to the original non-contrastive setting. IA methods in this setting measure how much an input token $x_{i}$ influences the model to increase the probability of target token $y_{t}$ while decreasing that of foil token $y_{f}$. A foil token can be defined as an output with an alternative answer, suffering from incorrect generalization (§ 3). Contrastive versions of IE, GN, and I×$\times$×G are defined as follows:
@@ 2.2 Interpreting in-context learning (ICL)
We focus on the ICL setting Brown et al. (2020), which has typically been adopted in modern LLM-based reasoning. An input prompt in ICL setting consists of few-shot examples $E$ and a target question. $E$ is composed of $n$ examples [$e_{1},\cdots,e_{n}$], each of which contains an input-output pair $e_{i}=(X_{i},f(X_{i}))$, given a function $f$ associated with the task. Let $X_{n+1}$ represent the target question $q$ that the model must answer. The ICL setting is denoted as follows:
$\underbrace{\overbrace{\overbrace{X_{1},f(X_{1})}^{\textit{example}~{}e_{1}},\ldots,\overbrace{X_{n},f(X_{n})}^{\textit{example}~{}e_{n}}}^{\textit{few-shot examples E}},\overbrace{X_{n+1}}^{q}}_{\textit{prompt}\ X},\underbrace{f(X_{n+1})}_{\textit{completion\ y}}$
@@
Notably, we consider the use of IA methods to identify a particular example $e^{*}\in E$ in input, instead of a particular token. To compute an IA score for an example $S(e_{i})$, we sum up the IA scores for its constituent tokens [^2] See Appendix A for results obtained with max aggregation instead of sum; see Appendix B for results that treat the tokens in each example as a coalition and compute their attribution.: $S(e_{i})=\sum_{x_{j}\in(X_{i},y_{i})=e_{i}}S(x_{j})$. [^3] An exception applies in the IE method; the attribution score for an example $e_{i}$ is simply computed by erasing the corresponding $X_{i}$ and $f(X_{i})$ from the input sequence. Our interest is which example obtains a high IA score.
@@ 3 Problem settings
Figure 2: IA scores (attribution accuracies) for each task/model in the 10-shot setting (thus, the chance rate is 20% and 10% for the top-two and top-one metric, respectively; red dotted line). The edit distance and attention baselines are indicated by a black dotted line and gray bar, respectively.
In addition to them, we adopted a simple task of associative recall (AR), which is typically employed in studying ICL, where the model is supposed to simply memorize the key: value mapping rules demonstrated in the prompt and apply them to the target question. Linguists may be more interested in the task of, for example, syntactic transformation to an interrogative sentence McCoy et al. (2020) based on the original poverty of the stimulus argument in the language domain Chomsky (1980). However, the models’ inference on such a realistic task may interfere with their meta-linguistic knowledge, which is inherently hard to track with IA methods; thus, we adopted synthetic tasks. [^4] See Appendix F for results on tasks with chain-of-thought format.
@@ 4.1 Overview
For each synthetic task, we create 360 different questions with different sets of few-shot examples and a target question. In the LD, AM, VO, TA, and PT tasks, the correct rule is selected out of the two candidates (rules A or B shown in Table 1) in a 1:1 ratio. The position of the disambiguating example $e^{*}$ is randomly selected according to a uniform distribution over all positions except the first.
@@
We evaluate six LLMs: Llama-2-7B, Llama-2-13B Touvron et al. (2023), Gemma-2-2B, Gemma-2-9B, Gemma-2-27B Riviere et al. (2024), and Mistral-7B Jiang et al. (2023). As a prerequisite for our experiments, the models should be able to learn and solve the task, i.e., be sufficiently sensitive to the disambiguating example $e^{*}$ and use this to determine the correct rule. To ensure this ability, we fine-tune these models on each task (see Appendix C), but the conclusions overall did not alter before and after fine-tuning (see Appendix D). Furthermore, we excluded the instances where models yielded an incorrect answer, as interpreting the model’s failure is beyond our scope.
@@ 4.3 Baselines
We also examine directly asking the models to generate their rationale. Specifically, we have models generate the most informative example in a prompt (Appendix E) in deriving their answer as a post-hoc explanation. This might be, more or less, relevant to the verbalization of aha moment recently observed in DeepSeek models Guo et al. (2025).
@@ 5 Experiments
Figure 2 shows the results in 10-shot settings, with both top-2 and top-1 metrics. Additional analyses are presented in Appendices AF, and G.
@@ 5.2 Scaling properties
Figure 3: IA score (attribution accuracy) for interpreting Gemma-2-2B models across all six tasks. Gradient-based methods are relatively robust to the number of few-shot examples, while there is a consistent, large drop in attribution accuracy in SA. Note that both x-axis and y-axis are in log scale.
@@
Next, given the trend of long-context LLMs, we examine the relationship between attribution accuracy and the number of few-shot examples. Figure 3 shows the attribution accuracy for Gemma-2-2B in all six tasks in different numbers of in-context examples. This demonstrates that gradient-based methods maintain accuracy or rather improve against the longer context, in contrast to the decreasing chance rate. This suggests the robustness of IA methods in long-context scenarios, highlighting their potential for interpreting inputs with extensive contextual information. Notably, the quality of self-answer consistently degraded as the number of in-context examples increased, while the positive scaling effect was observed in the previous analysis of model size. That is, the gradient-based methods and self-answer approach exhibit an insightful trade-off between different scaling properties.
@@
TA (Rule A) $98.0$ $52.5$ $50.0$
TA (Rule B) $2.0$
PT (Rule A) $68.5$ $59.0$ $51.1$
PT (Rule B) $31.5$
@@ 6 Discussion
Our tasks can not be answered without disambiguating aha examples. Thus, it is somewhat unintuitive to see the non-100% attribution accuracy of the IE method (again, the LLMs understood the task as they achieved 100% accuracy in the tasks) — what happened here? To obtain a hint to clarify IE’s potential limitations, we analyze model behaviors when the disambiguating example is excluded. Interestingly, LLMs adopted a specific generalization (rule) in each task when there was no disambiguating example (Table 2); in other words, they sometimes exhibited strong inductive biases in our tasks. That is, when the correct rule is equal to their preferred rule by their inductive bias, they can answer the task correctly even without disambiguating examples, and the IE method does not compute a proper attribution score. It is now common to see LLMs have particular inductive biases (not a tabula rasaWarstadt et al. (2020b); Kharitonov and Chaabouni (2021). Catching such generalization bias with IA methods represents an inherent challenge, highlighting their potential limitations in interpreting LLMs.
Why was I×$\times$×G worse than GN?
@@
The advantage of the I×$\times$×G method, compared to the GN, is the consideration of the norm of the input embedding Shrikumar et al. (2017); Denil et al. (2014). Since a large vector tends to have a large dot product with another vector, the norm of the input token embedding (vector) is expected to affect the IA score of I×$\times$×G. Then, no improvement of the I×$\times$×G over the GN suggests that, at least in our settings, the norm of the embeddings was not informative to estimate the input attribution. The norm of the embedding largely has decontextualized information about the word, such as frequency, and it may make sense that such information is not helpful to interpreting our controlled, synthetic ICL tasks consisting of alphabet characters, numbers, or random words.
Why was IG worse than GN?
@@
IG is a path-based approach; the gradient is accumulated from a baseline vector (typically a zero vector) to the targeted input representation (in our case, the sequence of input embeddings representing few-shot examples). This approach is somewhat intuitive when considering an attribution for a particular word or sentence; for example, suppose one computes an attribution to the word “excellent” in an input, IG may trace the path from zero to the “excellent” vector, which will go through the goodness direction, involving the intermediate points corresponding to, e.g., “okay” “decent,” “good,” “excellent” Sanyal and Ren (2021). Then, one critical question is — what does this path mean in interpreting the prompt and task representations? Different prompt representations will no longer correspond to the same targeted task; thus, the attribution of a particular token under the intermediate point to the target prompt vectors may no longer be an attribution under a targeted task. This can be one concern regarding the ineffectiveness of IG in our settings.
Figure 4: IA score (attribution accuracy) obtained with different baselines for IG when interpreting the Gemma-2-2B models across all six tasks. Zero vectors consistently yield higher accuracy than <pad> tokens.
Figure 5: IA score (attribution accuracy) for different disambiguating, aha example positions. Gemma-2-2B model is used. All methods exhibit some degree of positional bias toward the aha example’s position.
Will the Baseline Choice for IG Change the Results?
Existing work on BERT Devlin et al. (2019) shows that the [MASK] token is a better baseline for IG, whereas the [PAD] token and zero vectors yield no substantial difference in performance Bastings et al. (2022). Although we adopted zero vectors as the baseline in our main experiments, given that causal LMs do not have mask tokens, we further discuss the impact of this baseline choice on our results. We conduct a case study on the Gemma-2-2B models, using pad tokens (<pad>) as an alternative baseline for IG. The results are shown in Figure 4. The IA scores obtained with zero vectors consistently surpass those obtained with <pad> tokens, suggesting that zero vectors are the better choice for interpreting causal LLMs.
Positional Impact of the Aha Example
LLMs are known to exhibit positional bias to prioritize information at the beginning or the end of the input Liu et al. (2024). Given this concern, we also examine the positional bias underlying our interpretability results. Figure 5 shows attribution accuracy by the position of disambiguating, aha example. IA methods exhibited positional bias, though the patterns varied across tasks and methods. For instance, in the VO task, IE and GN performed better when the aha example appeared at the beginning, while in the AR task, IG performed better when it appeared at the end. See Appendix G for other models’ results as well as for 50 and 100-shot settings.